# Kashmir-190006, India

### v3

@@ 1 Introduction
The apparent loss of unitarity during black hole evaporation (aka the black hole information paradox) has remained one of the most profound challenges in attempting to reconcile quantum mechanics with gravity. A crucial insight Page:1993df was that for black hole evaporation to be unitary, the entanglement entropy of radiation should initially increase but eventually decrease, following what is now known as the Page curve. It was only recently shown that at least for certain black holes, the entanglement entropy during evaporation indeed follows the Page curveAlmheiri:2019hni ; Penington:2019npb .
@@
In Almheiri:2019hni , for example, the set-up considered was a two-dimensional black hole in Jackiw-Teitelboim gravity attached to a bath. By computing the quantum extremal surface via the generalized entropy formula of Engelhardt and WallEngelhardt:2014gca , it was found that there are two extremal surfaces corresponding to two saddles of the gravitational path integral—the Hawking saddle and the island saddle. A switch between the saddles takes place at Page time, and the Page curve is recovered. Since these breakthroughs, a deeper understanding of the transition has been obtained in terms of replica wormholes Penington:2019kki ; Almheiri:2019qdq ; Almheiri:2020cfm . Building on these developments, there have been many investigations of Page curves for various black holes Almheiri:2019yqk ; Marolf:2020xie ; Almheiri:2019psy ; Chen:2020uac ; Hashimoto:2020cas ; Hartman:2020khs ; Gautason:2020tmk ; Balasubramanian:2020xqf ; Krishnan:2020fer ; Krishnan:2020oun ; Ling:2020laa ; Ahn:2021chg .
In general, the evaporation of a black hole will be influenced by external conditions, such as ambient temperature or energy exchange with other systems. It is interesting to ask how changing the external conditions affects the Page curve? In this paper, we address this question in the same set-up involving a black hole coupled to a bath used in Almheiri:2019hni , where we model the changing external conditions by changing the bath temperature.
We consider a bath whose temperature continuously evolves with time via some given formula. As the temperature of the bath changes, it leads to a change in the boundary parameterization $f(t)$ that describes the black hole geometry in JT gravity. This in turn modifies the positions of the quantum extremal surface (QES) at different times, and hence the Page curve is also modified. We compute the modified Page curves numerically.
@@
In more detail, we consider three different cooling profiles– instantaneous temperature drop(for the bath), linear drop in energy, and the temperature following Newton’s law of cooling. For each case, we track the subsequent evolution of the boundary parameterization numerically, starting from the parameterization corresponding to a black hole at temperature $\beta$ and reaching a temperature $\tilde{\beta}$. Considering the process to be slow enough to remain within the semi-classical regime, we compute the positions of the quantum extremal surfaces and the corresponding entropies at each time step during the process of settling into a new temperature. The results are the Page curves of evaporating black holes in their non-equilibrium phases for each of the three cooling profiles. Note that while we have restricted ourselves to these profiles, the numerical method we use can be utilized for any general case as long as it stays in the semiclassical regime.
A scenario in a somewhat similar spirit was explored in Hollowood:2020cou . This paper considered the modification of the Page curve due to shock waves sent from the bath to the black hole. In this paper, which obtained analytical results, the increase of temperature was assumed to be instantaneous. Our investigations are numerical, which allows us to go beyond these limiting assumptions and investigate more general cases where the temperature of the bath is changed continuously with time.
@@
The structure and outline of the paper are set out as follows. In Section II, we start with preliminaries and provide the setup. Section III presents the computations for the instantaneous drop in temperature. In section IV, we generalise to more general cooling profiles. Section V. We conclude with a brief summary of the new results.
@@ 2.1 Jackiw-Teitelboim Gravity
Here, we outline the basic aspects of Jackiw-Teitelboim(JT) gravityJackiw:1984je ; Teitelboim:1983ux and nearly AdS2 holographyMaldacena:2016hyu ; Almheiri:2014cka ; Jensen:2016pah ; Maldacena:2016upp ; Engelsoy:2016xyb . We refer the reader to Mertens:2022irh for a recent detailed review.
@@
Here $\phi$ denotes the dilatonic field. The gravitational action is given by:
@@ 2.2 Quantum Extremal Surfaces
The entropy of Hawking radiation and the generalised entropy are given respectively by the formulae:
@@
Here, $G_{N}$ denotes Newton’s constant, $R$ is the radiation region, $I$ is the island, $\partial I$ is the boundary of the island, and $S_{\text{QFT}}(R\cup I)$ represents the entanglement entropy of the quantum fields in the bulk region of spacetime. The term $R\cup I$ denotes the union of the island and the radiation region. Extremization of $S_{\text{gen}}$ determines the location of the island. The surfaces that extremize the generalised entropy are called quantum extremal surfaces (QES).
@@
When there is a non-trivial QES in the bulk, it corresponds to an island. Figure 1 shows a QES where the exact position of ‘a’ is determined by extremization.
@@ 2.3 Black Hole at Equilibrium
$u=\frac{x^{+}+x^{-}}{2}\in\mathbb{R},\quad s=\frac{x^{+}-x^{-}}{2}>{\epsilon}$
The AdS2 boundary is parametrized by a coordinate transformation $u=f(t)$ which relates the Poincaré coordinates $x^{\pm}$ to the Rindler coordinates $y^{\pm}$. The metric in the $y^{\pm}$ coordinates is given by:
@@
We attach a half-Minkowski thermal bath at the holographic time-like boundary at $z=-\epsilon$. Note that this serves as our IR cut-off. We use flat coordinates for the bath, and we do so by analytically continuing the coordinates of the bulk.
@@
$x^{\pm}=\frac{\beta}{\pi}\tanh\left(\frac{\pi y^{\pm}}{\beta}\right)$ (10)
. Then the black hole patch has a metric of the form:
@@
The ADM energy of spacetime is holographically related to the boundary parameterization:
@@
$\{f(t),t\}$ is the Schwarzian derivative of the boundary parametrization. At the boundary, the rate of change of energy is:
@@
Where $T_{y^{+}y^{+}}$ is the ingoing stress-energy tensor that originates from the bath attached to the black hole, and $T_{y^{-}y^{-}}$ is the outgoing stress-energy tensor, i.e., the Hawking radiation from the black hole.
@@
The stress-energy tensor components transform as:
[…]
$T_{x^{\pm}x^{\pm}}\right>_{\eta}=\left(\frac{\partial y^{\pm}}{\partial x^{\pm}}\right)^{2}\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}+\frac{c}{24\pi}\{x^{\pm},y^{\pm}\}$ (17)
Equation (15) describes a Weyl transformation of the stress-energy tensors, (16) describes the simple tensorial transformation, while (17) represents a coordinate transformation accompanied by a Weyl transformation, respectively.
In the case of an eternal black hole in thermal equilibrium with the radiation bath, both the incoming and outgoing modes of the stress tensor have temperature $\frac{1}{\beta}$
$T_{y^{\pm}y^{\pm}}\right>=\frac{\pi c}{12\beta^{2}},$ (18)
Note that these are expectation values of the normal ordered stress tensor with respect to a flat metric in the respective coordinate.
@@
We can transform our value of stress tensor using (17), and in doing so, we get $T_{x^{\pm}x^{\pm}}\right>=0$ as expected
2.4 Non-Equilibrium
In the previous section, we described a black hole at equilibrium with a heat bath. In this section, we will consider a non-equilibrium situation. In order to create an evaporating black hole, the temperature of the black hole must be higher than that of the bath. In this scenario, radiation can leak into the bath, leading to the evaporation of the black hole.
The change in the bulk geometry because of this is captured in the boundary parameterization in the semi-classical regime, where we assume that there is no back reaction on the geometry.
We start with the bath and the black hole at equilibrium at inverse temperature $\beta$ and then make a sudden change at boundary time $t=0$. Before the change, at $t<0$ or alternately $y^{+}<0$, we have:
$T_{y^{+}y^{+}}=\frac{\pi c}{12\beta^{2}}$ (20)
Then suddenly the temperature of the bath changes to $\tilde{\beta}$ at the boundary tie $t=0$. This change propagates to the bulk causally. After $y^{+}>0$, we get :
$T_{y^{+}y^{+}}=\frac{\pi c}{12\tilde{\beta}^{2}}$ (21)
$T_{y^{-}y^{-}}=-\frac{c}{24\pi}\{f(y^{-}),y^{-}\}.$ (22)
Energy conservation at the holographic boundary gives from (14):
Upon substituting the parameterization with E, we get:
$E=e^{-kt}\int\left[\frac{\pi c}{12\tilde{\beta^{2}}}e^{kt}dt\right]$ (25)
Where $k=\frac{G_{N}c}{3\phi_{r}}$.
We solve this differential equation with the boundary condition that at $t=0$ the ADM energy is $E=\frac{\pi\phi_{r}}{4\beta^{2}G_{N}}$, which was the original energy of the black hole before any evaporation took place. The third-order differential equation for boundary parametrization is obtained as:
$\{f,t\}=-2\pi^{2}[\tilde{\beta}^{-2}-(\tilde{\beta}^{-2}-\beta^{-2})e^{-kt}].$
This is solved in Almheiri:2019qdq Hollowood:2020cou where the solution for boundary parameterization comes out to be modified Bessel functions. Note we fix $k<<1$ in order to work in the semi-classical regime. In our case, $k=0.05$, and also the time step for evolution is 0.01.
However, we generally proceed with the differential equation and solve it numerically. This numerical solution allows us the flexibility to solve some other variations of this differential equation that are not solvable by analytical methods. Our result for a single sudden change in temperature of the bath is analogous to the result Hollowood:2020cou where the black hole evaporates from one temperature and settles down to another. Although the physical process remains the same in both cases, our numerical approach allows for both the $w\pm$ coordinates to change and fine-tuning of the temperature change profile, whereas the result Hollowood:2020cou is analytical in nature and for a single instant change in temperature of the black hole using a shockwave. This keeps the bath temperature constant throughout the analysis, thereby only changing the outgoing ‘w’ coordinate.
2.5 Early Times and Late Times
In early times, the black hole was in equilibrium with the bath at temperature $T_{\beta}$, and we are interested in black hole evaporation at very late times when the black hole has reached equilibrium with the bath at temperature $T_{\tilde{\beta}}$. Solving the differential equation numerically with the appropriate boundary condition results in interpolation between the initial and final parameterization, which is what we expected.
Initially:
At very late times, it should approximate:
Figure 2: Plot of first derivative of $f(u)$ with $u$
Figure 3: Plot of second derivative of $f(u)$ with $u$
From Figure 3 and 3, it can be seen how the derivative and double derivative of the boundary parameterization interpolate between the old and very late times (new) boundary parameterizations.
3 Evaporation of black hole: instantaneous change in bath temperature
In this section, we present our computations for black hole evaporation when the temperature of the bath drops instantaneously. The evaporation of the black hole and its dynamics are captured in the parameterization, which interpolates between the old and new parameterizations. We track the evolution of the entropy of the black hole using (5) and obtain the Page curve of an evaporating black hole. The entropy is calculated numerically for the Hawking Saddle and the Island Saddle. We evaporate the black hole by decreasing the temperature of the bath and breaking the equilibrium.
@@ 3 Evaporation of black hole: instantaneous change in bath temperature
In this frame, the matter entanglement entropy between two points $a$ and $b$ is given by the Cardy-Calabrese formula Calabrese:2004eu :
@@
Here, ${\epsilon^{w^{+}_{a}},\epsilon^{w^{-}_{a}},\epsilon^{w^{+}_{b}},\epsilon^{w^{-}_{b}}}$ are the UV cutoffs for the entropy at points $a$ and $b$. The cut-offs are taken with respect to the proper distance in the physical metric. Saddles.
Now we want to relate the $w-$frame to the original coordinate system $y$. The stress energy tensor in the two frames is related by:
@@
Likewise, from (21):
$\left\{w^{+},y^{+}\right\}=-\frac{2\pi^{2}}{\widetilde{\beta}^{2}},\quad y^{+}\geqslant 0$ (38)
@@
As (38) is a third-order differential equation, we have to specify three initial conditions at $y^{+}=0$. We impose continuity in the $w^{+}$ coordinates up to the second-order derivative. This gives us:
$w^{+}|_{y^{+}=0}$ $=1$
@@
The discontinuous change in the bath temperature is reflected as the discontinuity of the Schwarzian, which shows up in the third-order derivative. This discontinuity is an artifact of our rather artificial assumption of a discontinuous change in the bath temperature.
@@
$\left\langle T_{y^{-}y^{-}}\right\rangle=\frac{\pi c}{12\beta^{2}},\quad y^{-}<0,$ $\left\langle T_{y^{-}y^{-}}\right\rangle=\frac{\pi c}{12}\left[\frac{1}{\tilde{\beta}^{2}}-\left(\frac{1}{\tilde{\beta}^{2}}-\frac{1}{\beta^{2}}\right)e^{-kt}\right]\quad y^{-}\geqslant 0$ (39)
It follows that for $y^{-}<0$, the relation between $y^{-}$ and $w^{-}$ has the same form as (36):
@@
Which is solved by $w^{-}=-e^{\frac{-2\pi}{\beta}}$.
@@
$\left\{w^{-},y^{-}\right\}=$ $-2\pi^{2}\left[\tilde{\beta}^{-2}-\left(\tilde{\beta}^{-2}-\beta^{-2}\right)e^{-kt}\right],\quad y^{-}\geqslant 0$ (41)
@@
Unlike in the previous case, here the Schwarzian is no longer discontinuous. This reflects the fact that although the bath temperature changes suddenly, the black hole temperature changes in a continuous manner.
@@
Where $f$ is the boundary parameterization function. Note that for the point $a$ in the bulk, $\Omega^{2}$ varies, but for the point $b$ in the bath, it is a constant.
@@
Now we are ready to calculate the entropy using the Island formulae (5) and (6).
3.1 Hawking saddle
We start with the calculation of the trivial Hawking saddle.
[[[\,[$b_{L}$]]\,]]$b_{R}$[[[\,[$b_{L}$]]\,]]$b_{R}$

### v4

@@ 1 Introduction
The evaporation of black holes has been a topic of fascination for physicists for many years. According to the principle of unitarity, the entanglement entropy of radiation should initially increase but eventually decrease during the evaporation of a black hole, following what is now known as the Page curve Page:1993df. Obtaining a Page curve for black hole evaporation remained an open challenge for a long time. Only recently was it shown that for certain black holes, the entanglement entropy during evaporation indeed follows the Page curve Almheiri_2019; Penington:2019npb; Almheiri:2019hni.
@@
In Almheiri:2019hni, for example, the set-up considered was a two-dimensional black hole in Jackiw-Teitelboim gravity attached to a bath. By computing the quantum extremal surfaces via the generalized entropy formula of Engelhardt and WallEngelhardt:2014gca, the Page curve was recovered. Since then, there has been significant progress in understanding the results and extending them to various black holes Penington:2019kki; Almheiri:2019qdq; Almheiri:2020cfm; Almheiri:2019yqk; Marolf:2020xie; Almheiri:2019psy; Chen:2020uac; Hashimoto:2020cas; Hartman:2020khs; Gautason:2020tmk; Balasubramanian:2020xqf; Krishnan:2020fer; Krishnan:2020oun; Ling:2020laa; Ahn:2021chg.
Most work on this topic has focused on black holes interacting with a bath that remains at a fixed temperature. However, the evaporation of a black hole can be influenced by external conditions such as changes in ambient temperature or energy exchange with other systems. For instance, primordial black holes in the early universe evaporate in an environment with rapidly changing ambient temperature. A natural question to ask is: how does changing the external conditions in a time-dependent manner affect the Page curve? In other words, how robust is the Page curve under time-dependent externalities?
In this paper, we address this question in the same setup considered in Almheiri:2019hni: a 2D black hole coupled to a reservoir. In our case, the reservoir is taken to be a cooling bath: one with controllable, time-dependent temperature. This provides a simple toy model for non-equilibrium evaporation.
The results from this model carry over beyond the case of a cooling bath. They apply to all processes where the black hole loses (or gains) mass in a time-dependent manner. Examples include the injection of negative mass shells into the black hole Cadoni:2021ypx and partially reflecting boundary conditions with time-dependent reflectivity. As we will show, all such setups for a black hole out of equilibrium with the bath are dependent on a single parameter, and hence mathematically equivalent.
In JT gravity, the black hole geometry is described by boundary parametrization $f(t)$. As the temperature of the bath changes, it leads to a change in $f(t)$. This in turn shifts the location of the quantum extremal surface (QES), thereby modifying the Page curve. This is the modification that we aim to study in this paper. There are a few special situations where analytic control is possible, such as instantaneous temperature jumps sourced by shock waves from the bath to the black hole Hollowood:2020cou. Outside of these cases, the evolution of $f(t)$, and therefore the Page curve, can not be computed analytically.
We therefore adopt a numerical approach that allows for arbitrary profiles beyond analytic solvability. We consider scenarios where the bath temperature is allowed to change continuously with time and compute the modified Page curves numerically for various cooling profiles. In the limiting case of shock waves, we demonstrate that our results are in agreement with those of Hollowood:2020cou, where they were computed analytically.
In more detail, we consider three different cooling profiles. They are: an instantaneous temperature drop (for the bath), linear drop in energy, and the temperature following Newton’s law of cooling. These profiles are chosen to model environments where a black hole exchanges energy with an external system whose properties evolve in a controllable way. Our only assumption is that the process is slow enough to remain within the semiclassical regime (back-reaction remains negligible).
@@
For each case, we track the subsequent evolution of the boundary parametrization numerically utilizing the RK4 method, starting from the parametrization corresponding to a black hole at temperature $\beta$ and reaching a temperature $\tilde{\beta}$. We compute the positions of the quantum extremal surfaces and the corresponding entropies at each time step during the relaxation to a new temperature. The results are the Page curves of evaporating black holes in their non-equilibrium phases for each of the three cooling profiles.
Note that while we have restricted ourselves to these profiles, the numerical method we use can be utilized for any general case as long as it stays in the semiclassical regime.
To summarize, our main new contribution is extending the computation of the Page curve to cases where the black hole is interacting with a controlled, time-dependent environment (which changes the boundary parametrization in a continuous manner) utilizing numerical methods.
@@
The structure and outline of the paper are set out as follows. In Section II, we start with preliminaries and provide the setup. Section III presents the computations for the instantaneous drop in temperature. In Section IV, we generalize to more general cooling profiles. We conclude with a brief summary of the new results and future directions.
@@ 2.1 Jackiw-Teitelboim Gravity
We start by reviewing the basic aspects of Jackiw-Teitelboim(JT) gravity Jackiw:1984je; Teitelboim:1983ux and nearly AdS2 holography Maldacena:2016hyu; Almheiri:2014cka; Jensen:2016pah; Maldacena:2016upp; Engelsoy:2016xyb. We refer the reader to Mertens:2022irh for a recent detailed review.
@@
Here, $\Phi$ denotes the dilaton field. The gravitational action is given by:
@@ 2.2 Quantum Extremal Surfaces
The entropy of Hawking radiation and the generalized entropy are given, respectively, by the following formulae:
@@
Here, $G_{N}$ denotes Newton’s constant, $R$ is the radiation region, $I$ is the island, $\partial I$ is the boundary of the island, and $S_{\text{QFT}}(R\cup I)$ represents the entanglement entropy of the quantum fields in the bulk region of spacetime. The term $R\cup I$ denotes the union of the island and the radiation region. Extremization of $S_{\text{gen}}$ determines the location of the island. The surfaces that extremize the generalized entropy are called quantum extremal surfaces (QES).
@@
When there is a non-trivial QES in the bulk, it corresponds to an island. Figure 1 shows a QES where the exact position of ‘a’ is determined by extremizing the generalized entropy.
@@ 2.3 Black Hole at Equilibrium
$u=\frac{x^{+}+x^{-}}{2}\in\mathbb{R}\quad\text{and}\quad s=\frac{x^{+}-x^{-}}{2}>{\epsilon}$
The AdS2 boundary is parametrized by a coordinate transformation $u=f(t)$ that relates the Poincaré coordinates $x^{\pm}$ to the Rindler coordinates $y^{\pm}$. The metric in the $y^{\pm}$ coordinates is given by:
@@
We attach a half-Minkowski thermal bath at the holographic time-like boundary at $z=-\epsilon$. Note that this serves as our IR cutoff. We use flat coordinates for the bath and we do so by analytically continuing the coordinates of the bulk.
@@
$x^{\pm}=\frac{\beta}{\pi}\tanh\left(\frac{\pi y^{\pm}}{\beta}\right).$ (10)
Then the black hole patch has a metric of the form:
@@
The ADM energy of spacetime is holographically related to the boundary parametrization:
@@
$\{f(t),t\}$ is the Schwarzian derivative of boundary parametrization. At the boundary, the rate of change of energy is:
@@
Here $T_{y^{+}y^{+}}$ is the ingoing stress-energy tensor that originates from the bath attached to the black hole, and $T_{y^{-}y^{-}}$ is the outgoing stress-energy tensor, i.e., the Hawking radiation from the black hole.
@@
The stress-energy tensor components transform as follows:
$\left<T_{y^{\pm}y^{\pm}}\right>_{g}=\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}-\frac{c}{12\pi}\frac{\partial_{\pm}^{2}\Omega}{\Omega},$ (15)
$\left(\frac{\partial x^{\pm}}{\partial y^{\pm}}\right)^{2}\left<T_{x^{\pm}x^{\pm}}\right>_{\eta}=\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}+\frac{c}{24}\{x^{\pm},y^{\pm}\}.$ (16)
Equation (15) describes a Weyl transformation of stress-energy tensors, here $\left<T_{y^{\pm}y^{\pm}}\right>_{g}$ corresponds to the normal ordered stress tensor with respect to the frame with metric $ds^{2}=-\Omega^{2}(y^{+},y^{-})\,dy^{+}dy^{-}$, and $\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}$ corresponds to the normal ordered stress tensor with respect to the frame with metric $ds^{2}=-dy^{+}dy^{-}$. Equation (16) represents a complete conformal transformation, where, $\left<T_{x^{\pm}x^{\pm}}\right>_{\eta}$ corresponds to the normal ordered stress tensor with respect to the flat metric $ds^{2}=-dx^{+}dx^{-}$, and $\left<T_{y^{\pm}y^{\pm}}\right>_{\eta}$ corresponds to the normal ordered stress tensor with respect to the flat metric $ds^{2}=-dy^{+}dy^{-}$ .
In the case of an eternal black hole in thermal equilibrium with the radiation bath, both the incoming and the outgoing modes of the stress tensor have temperature $\frac{1}{\beta}$
$\left<T_{y^{\pm}y^{\pm}}\right>=\frac{\pi c}{12\beta^{2}},$ (17)
Note that the expectation values from this point onward are normal ordered stress tensor with respect to a flat metric in the respective coordinate. We remove the subscript from the further discussion.
@@
We can transform our stress tensor value using (16), and in doing so we get $\left<T_{x^{\pm}x^{\pm}}\right>=0$ as expected.
2.4 Entropy in terms of boundary parametrization
@@
In this frame, the matter contribution to entanglement entropy between two points $a$ and $b$ is given by the Cardy-Calabrese formula Calabrese:2004eu:
@@
Here, $\epsilon^{w^{+}_{a}},\epsilon^{w^{-}_{a}},\epsilon^{w^{+}_{b}},\epsilon^{w^{-}_{b}}$} are the UV cutoffs for the entropy at points $a$ and $b$. The cutoffs are taken with respect to the proper distance in the physical metric.
Now we want to relate the $w\pmframe$ to the original coordinate system $y\pm$. The stress energy tensor in the two frames is related by:
@@
Similarly, from equation (39):
$\left\{w^{+},y^{+}\right\}=-\frac{2\pi^{2}}{{\beta_{bath}}^{2}},\quad y^{+}\geqslant 0$ (25)
@@
As (25) is a third-order differential equation, we have to specify three initial conditions at $y^{+}=0$. We impose continuity in the $w^{+}$ coordinates up to the second-order derivative.
This choice comes from the fact that a discontinuity in $w,w^{\prime}$ or $w^{\prime\prime}$ would correspond to unphysical discontinuities in the geometry, while the change (may be discontinuous) in the third derivative is precisely what encodes the change (may be instantaneous) in energy flux from the bath.
This matches the physical expectation that a sudden temperature change should not introduce discontinuities in the geometry. Rather, it should produce a finite change in the stress tensor flux.
This gives us:
$w^{+}|_{y^{+}=0}$ $=1$ (26)
@@
For the case of discontinuous change in the bath temperature, the discontinuity is reflected as the discontinuity of the Schwarzian in the third-order derivative. This discontinuity is an artifact of our rather artificial assumption of a discontinuous change in the bath temperature.
@@
$\left\langle T_{y^{-}y^{-}}\right\rangle=\frac{\pi c}{12\beta^{2}},\quad y^{-}<0,$ $\left\langle T_{y^{-}y^{-}}\right\rangle=g(t)\geqslant 0$ (27)
Here, $g(t)$ is obtained by solving (14) for a specific bath cooling profile. It follows that for $y^{-}<0$, the relation between $y^{-}$ and $w^{-}$ has the same form as (23):
@@
This is solved by $w^{-}=-e^{\frac{-2\pi}{\beta}}$.
@@
$\left\{w^{-},y^{-}\right\}=$ $g(y^{-})\geqslant 0$ (29)
@@
Unlike in the previous case, here the Schwarzian is never discontinuous. This reflects the fact that although the bath temperature may change suddenly, the black hole temperature changes in a continuous manner.
@@
Where, $f$ is the boundary parametrization function. Note that for the point $a$ in the bulk, $\Omega^{2}$ varies, but for the point $b$ in the bath, it is a constant.
@@
Now we are ready to calculate the entropy of the Hawking saddle and Island saddle.
3 Dynamical Setup and Numerical Framework
In this section, we will set up the non-equilibrium problem of a black hole interacting with a bath of changing temperature. We will introduce our numerical methods and use them to obtain the Page curve for the case of an instantaneous change in temperature. We will see that our numerical results match with the analytical results obtained by Hollowood:2020cou to a very high degree of accuracy. This serves as a check of our numerical methods.
3.1 Breaking equilibrium: time-dependent bath temperature
The equilibrium in (14) is broken when the incoming stress tensor bath is modified in any manner, leading to the following:
$\frac{\partial E}{\partial t}=T_{y^{+}y^{+}}^{\text{modified}}-T_{y^{-}y^{-}}$ (37)
The energy of the black hole can be described as a function of the boundary parametrization. By plugging the incoming stress tensor, this boundary parametrization can be computed, and hence the entire evolution of the system can be tracked. Keeping all their parameters constant in (37), the dynamics of the black hole is completely captured by the time dependence of the incoming component of the stress tensor.
This implies that all situations where the stress tensor is modified in a time dependent way are mathematically equivalent. We study a cooling bath, but the method applies equally to for instance a situation where negative energy radiation is emitted from the black hole to the bath. Another example would be partially absorbing boundary condition. Since the evolution depends on a single parameter $T_{y^{+}y^{+}}^{\text{modified}}$, all of these different cases can be captured by (37).
The change in the bulk geometry because of this is captured in the boundary parametrization in the semi-classical regime. Note that we assume that there is no back reaction on the geometry.
3.2 Instantaneous change in bath temperature: Boundary parametrization
We now consider the case of an instantaneous change in the bath temperature. We start with the bath and the black hole in equilibrium at inverse temperature $\beta$ and then make a sudden change at boundary time $t=0$. Before the change, at $t<0$ or alternatively $y^{+}<0$, we have:
$T_{y^{+}y^{+}}=\frac{\pi c}{12\beta^{2}}$ (38)
The temperature of the bath instantaneously changes to $\tilde{\beta}$ at the boundary time $t=0$. This change propagates to the bulk causally. After $y^{+}>0$, we get :
$T_{y^{+}y^{+}}=\frac{\pi c}{12\tilde{\beta}^{2}}$ (39)
Note that this setup is mathematically completely equivalent to the one studied in Hollowood:2020cou, where a positive energy shock wave was sent to the black hole from the bath and the temperature of the black hole shifts and settles down to a new temperature. As a consistency check on our numerical methods, we will show that our results for boundary parametrization match with the analytical results of Hollowood:2020cou.
$T_{y^{-}y^{-}}=-\frac{c}{24\pi}\{f(y^{-}),y^{-}\}.$ (40)
Energy conservation at the holographic boundary given by (14):
Upon substituting the parametrization with E, we get:
$E=e^{-kt}\int\left[\frac{\pi c}{12\tilde{\beta^{2}}}e^{kt}dt\right]$ (43)
where $k=\frac{G_{N}c}{3\Phi_{r}}$.
We solve this differential equation with the boundary condition that at $t=0$ the ADM energy is $E=\frac{\pi\Phi_{r}}{4\beta^{2}G_{N}}$, which was the original energy of the black hole before any evaporation took place. The third order differential equation for boundary parametrization is obtained as:
$\{f,t\}=-2\pi^{2}[\tilde{\beta}^{-2}-(\tilde{\beta}^{-2}-\beta^{-2})e^{-kt}].$ (44)
At early times, the black hole was in equilibrium with the bath at temperature $T_{\beta}$, so:
At very late times, it once again settled to equilibrium at temperature $T_{\tilde{\beta}}$, so:
The initial conditions are chosen to ensure continuity of $f(t)$ up to the second derivative.
$f|_{t_{0}=0}$ $=0$ (51)
$\frac{df}{dt}\Big|_{t_{0}=0}$ $=1$
$\frac{d^{2}f}{dt}\Big|_{t_{0}=0}$ $=0$
Note that in general $\{f,t\}=g(t)$, where for the single instant drop in bath temperature gives the expression $g(t)=-2\pi^{2}\left[\tilde{\beta}^{-2}-\left(\tilde{\beta}^{-2}-\beta^{-2}\right)e^{-kt}\right]$. In general $g(t)$ would have to be solved from (14). The differential equation to solve for ‘f’ is the same as $w^{-}$(29), but with a slightly different initial condition as seen in (30) and (51). This makes them related by a Möbius transformation Hollowood:2020cou.
To ensure that the back-reaction on the geometry remains under control and that the semiclassical approximation is valid, we must have $k<<1$.
The exact solution of this differential equation was obtained in Almheiri:2019qdq; Hollowood:2020cou where $f(t)$ is expressible in terms of modified Bessel functions. The exact analytical solution for boundary parametrization is as follows:
$f(t)=\frac{\tilde{\beta}}{\pi}\frac{K_{\nu}(\nu z_{0})\left(\frac{\hat{f}(t)}{\hat{f}(t_{0})}-1\right)+z_{0}\tanh\frac{\pi t_{0}}{\tilde{\beta}}\left(\hat{f}(t)I^{\prime}_{\nu}(\nu z_{0})/\alpha-K^{\prime}_{\nu}(\nu z_{0})\right)}{K_{\nu}(\nu z_{0})\left(\frac{\hat{f}(t)}{\hat{f}(t_{0})}-1\right)\tanh\frac{\pi t_{0}}{\tilde{\beta}}+z_{0}\left(\hat{f}(t)I^{\prime}_{\nu}(\nu z_{0})/\alpha-K^{\prime}_{\nu}(\nu z_{0})\right)}$
where
$\hat{f}=\alpha\frac{K_{\nu}(\nu z)}{I_{\nu}(\nu z)},\qquad\nu=\frac{2\pi}{\tilde{\beta}k},\qquad z=\sqrt{\frac{E_{\text{shock}}}{E_{\beta}}}\,e^{-k(t-t_{0})/2}.$
The constant of normalization is fixed by using the initial condition on $\hat{f(t_{0})}=e^{2\pi t_{0}/\beta}$. Also for our case $t_{0}=0$. Here, the AdS radius is set to 1.
We now proceed to solve this equation numerically. The time step for evolution is taken to be 0.01. We use RK4 method for numerically solving the differential equation to obtain the boundary parametrization $f(t)$ and later for the $w\pm$’ coordinates.
First, we plot the first and second derivative of $f(t)$. This is shown in Fig 2(a) and Fig 2(b). As a consistency check, we match them with the analytically computed derivatives of $f(t)$ at early and late times.
(a) Plot of first derivative of $f(t)$ with $t$
(b) Plot of second derivative of $f(t)$ with $t$
Figure 2: Interpolation of derivatives
In the Figures 2(a) and 2(b), the numerically computed derivative is shown in red, while the analytically computed early and late time derivatives are shown in green and blue respectively. It can be seen how the derivative and double derivative of the boundary parametrization interpolate between the old and new boundary parametrizations.
(a) Plot of first derivative of $f(t)$ with $t$
(b) Plot of second derivative of $f(t)$ with $t$
Figure 3: Numerical check with analytical result
Figure 3(a) is a comparative plot between our numerical solution and the analytical solution of Hollowood:2020cou. The plots overlap entirely with only marginal accumulation error as shown in 3(b). This provides a check on our numerical methods.
RK4 typically has an error rate of the order $O(h^{4})$, where h (=0.01 in our case) is the time step. Our analysis is not very sensitive to this error; in fact, the standard Euler method for numerical integration also gives the same result. In order to satisfy the semiclassical regime, our k is set to “0.1” throughout the discussion. Other Parameters include $\beta=50\pi$ and $\tilde{\beta}=55\pi$, c=100 and $G_{n}=0.06$.
3.3 Instantaneous Change in Temperature: Entropy saddles
Now that we have computed the evolution of the boundary parametrization with time, we use it to compute the generalised entropy in the Hawking saddle and the Island saddle.
3.3.1 Hawking saddle
We start with the calculation of the Hawking saddle.
$b_{L}$$b_{R}$$b_{L}$$b_{R}$